# Reasoning Language Models: A Blueprint

### v3

@@ B.4.1 Evaluation Types
Absolute evaluations assign scalar values to each step, ssessing aspects such as coherence, correctness, or helpfulness, using regression-based models. Moreover, evaluation dimensions can also be modeled as binary with classification models. While regression models provide more information, classification models capture correctness more naturally since a statement is usually correct or incorrect. On the other hand, the former ones are more suitable for measuring quality, such as the degree of coherence. Depending on the specific quality being evaluated, the choice between regression and classification models should align with the evaluation’s goals. Additionally, absolute scores can be transformed into rankings if needed, providing flexibility across various applications.
@@ B.4.2 Evaluation of Reasoning Steps
External model approaches include value models that predict scalar reward signals (Reward models) [35, 151, 88], reinforcement learning values like state-values (V-value models) [134], state-action values (q-value models), or pairwise models like the Bradley-Terry and PairRM frameworks. A more detailed comparison of reward models, v-value, and q-value models can be found in Appendix B.3.2.
There exist a large range of internal model approaches as substitutes for value models. They typically rely on methods like prompting the policy to output scores. Examples include MCT Self-Refine (MCTSr) [176], querying for a binary feedback (e.g., “Is the answer correct? answer“yes” or “no””) [179] and evaluating the probability of the output, leveraging uncertainty metrics such as token entropy or aggregated probabilities [182], and others [178].
@@ C.1.1 Setup and Notation
We will now present the details of the training pipeline of x1.
MDP Design x1 assumes the MDP following the definition presented in Appendix A.1 with the $\gamma$ values between $[0.95,1]$ to avoid over-penalizing long reasoning sequences. In the RLM setup, the state space and action space of the underlying MDP constitute a tree in which every state $s$ other than the starting state $s_{0}$ has exactly one action $a_{s}$ leading to it. This allows us to simplify the notation by omitting actions wherever it’s clear from the context that we are referring to only action leading to a given. For every action $a$ leading from the state $s$ to the state $s^{\prime}$ we will write:
@@
$V_{\pi}(s_{t})$ $=\mathbb{E}_{\pi}\left[\gamma^{T-t}r(s_{T+1})\right]\quad\in[-1,1]$ (4)
@@
The policy model We use a pretrained LM with parameters $\theta$ as a policy model and denote it $\pi_{\theta}$. The model autoregressively generates a sequence of tokens. We use a special token ’End of Intermediate Step’ (eois) to indicate the end of the reasoning step. We use a standard end-of-sequence (eos) token to indicate the end of the final reasoning step concluding the reasoning trajectory.
The value model A parametric value model is used to evaluate the quality of states. While MCTS traditionally approximates these values through extensive simulations, such an approach is computationally expensive and impractical in the RLM context. Inspired by AlphaZero [134], which replaces simulations with a parameterized value model, we estimate state-action values (short q-value) for reasoning sequences using a value model — effectively employing a process-based q-value model $Q_{\varphi}$ (see Appendix B.3). The value model is instantiated as a pretrained transformer-based LM, modified by adding three linear layers and a shifted, rescaled sigmoid activation to align the output domain to the state action function domain $[-1,1]$ (see Eq. 5). This setup proved more stable than alternatives, such as a tanh activation or a cropped linear layer. We will show in the following how such a model can be trained and provide a description for the data generation process in Appendix D. During training, we assume access to a final answer verifier, which evaluates the correctness of the model’s final answer and provides the true reward.
@@ C.1.2 MCTS Algorithm
We now present the algorithmic steps of a Monte Carlo Tree Search variant similar to AlphaZero as implemented in the x1 reasoning framework. The MCTS search operates in two distinct modes: training and inference. The core difference is that, during training, a final answer verifier evaluates and scores the final reasoning steps, providing a true reward signal that is backpropagated through the MCTS tree. This reward serves as a reliable learning signal for the value model $Q_{\varphi}$. During inference, however, the verifier is unavailable, and decisions rely solely on the value model.
Notation. We chose to store all values in nodes instead of edges, which defines the following set of statistics saved for each node $s$:
  • $N(s)$ the visit count of node $s$
@@
  • $\beta(s)$ the binary terminality function, returns $1$ if the node $s$ is terminal 00 otherwise.
Selection. The selection phase iteratively identifies the most promising child node with a selection policy. We use the following selection policy which is the node-based variant of the PUCT algorithm in AlphaZero [135] (which is defined on edge-based values) without a prior for finding selecting a child of $s$:
$\operatorname*{arg\,max}_{s_{c}\in\mathcal{C}(s)}\quad q(s_{c})+\frac{\sqrt{N(s)-1}}{1+N(s_{c})}\cdot\left(c_{1}+\log{\frac{N(s)+c_{2}}{c_{2}}}\right)$ )
@@
Expansion. We append $M$ nodes to the selected leaf, $M$ being a hyperparameter. One of the major challenges in applying RLMs is maintaining the diversity of reasoning paths. By adding $M$ nodes, we increase the exploration of alternative reasoning paths.
Backpropagation. The backpropagation step serves to propagate information from the terminal nodes back to their ancestors. In our implementation, we update the running estimates of the q-values using the following formula:
@@
True Reward Propagation. We improve the quality of the q-values by propagating the real final rewards back through the tree when a terminal state $s_{T+1}$ is reached. During training, terminal nodes can be evaluated against a reference golden answer $g^{*}$ using an external verifier. For actions leading to terminal states, the associated reward is equal to the q-value see Eq. 5. Therefore, instead of using the prediction of the q-value model, we initialize $q(s_{T+1})$ with the true reward $r(s_{T+1})$ based on the evaluation of the external verifier. The reward is then backpropagated via the q-values through the tree with our backpropagation operator. This adjustment anchors the q-value model predictions with real reward signals and prevents the q-value model predictions to diverge.
Best Path Selection. After $N$ iterations, MCTS will have formed a tree in which every path corresponds to one of the explored reasoning trajectories. The final reasoning step in a path with the highest terminal value estimate is returned as the final solution.
Algorithm 1 MCTS for Reasoning (Training mode in blue)
@@
C.2 Training Phase 1
Overall Training Pipeline. To adequately employ the MCTS-based reasoning scheme introduced in the Appendix C.1, the policy model must be fine-tuned to generate responses in the format of semantically-relevant reasoning steps. The value model – a q-value model in our case – must be trained to accurately estimate the values of the sequences of reasoning steps.
Phase 1: Supervised Fine-Tuning. The first phase focuses on preparing the policy and value models to generate and evaluate reasoning trajectories effectively. This is achieved by supervised fine-tuning (SFT) training on a dataset of example sequences of reasoning steps (where intermediate reasoning steps are terminated by an ”End of Intermediate Step” eois token). The objective is twofold: (1) to fine-tune the policy model $\pi_{\theta}$ to produce semantically coherent reasoning steps, and (2) to train the q-value model $Q_{\varphi}$ to accurately assign scalar scores to reasoning trajectories, distinguishing between high-quality and suboptimal reasoning paths.
This supervised fine-tuning phase ensures that the policy can generate reasoning steps consistent with the structured format required for downstream MCTS-based exploration, while the q-value model provides reliable evaluations of intermediate and terminal states. Together, these components form the foundation for the subsequent online reinforcement learning in Phase 2, where the policy and q-value models are further refined through interaction with the reasoning framework.
C.2.1 Datasets Generation and Preparation
Dataset for SFT of the Policy. Performing SFT of the policy requires a dataset of high-quality reasoning sequences denoted as $D_{\text{SFT}}=\{\left(x_{\text{SFT}}^{(i)},\,y_{\text{SFT}}^{(i)}\right)\}$. Each pair in the dataset consists of a prompt $x_{\text{SFT}}^{(i)}$ composed of a sequence of reasoning steps (for example $x_{\text{SFT}}^{(i)}=(z_{0}^{(i)},...,z_{j}^{(i)})$), and a target completion $y_{\text{SFT}}^{(i)}=z_{j+1}^{(i)}$ which is the subsequent reasoning step or final answer. Appendix D contains a detailed account of the dataset creation and processing. It covers how the special eois token is appended to reasoning steps mark the end of a step during inference.
Dataset for Q-Value Model Training. Similarly to SFT, training the q-value model requires a supervised dataset of reasoning sequences and corresponding scores. We denote this dataset $D_{\text{QVM-train}}=\{(x_{\text{QVM-train}}^{(i)},y_{\text{QVM-train}}^{(i)})\}$, with reasoning sequences $x_{\text{QVM-train}}^{(i)}=(z_{0}^{(i)},...,z_{t}^{(i)})$ and target q-value $y_{\text{QVM-train}}^{(i)}$. Appendix D explains how this dataset can be generated using an initial list of questions, a base LLM for querying, and a verifier program to label reasoning sequences as conducive to a correct final answer or not.
@@ C.2.1 Datasets Generation and Preparation
Input: Policy Model $\pi_{\theta}$, tokenized dataset $D_{\text{SFT}}=\{(x^{(i)},y^{(i)})\}$, training hyperparameters (optimizer, learning rate $\eta$, batch size $B$, and maximum number of epochs $E$.).
@@
3:     Divide $D_{\text{SFT}}$ into batches $\{B_{k}\}$ of size $B$.
4:     for each batch $B_{k}$ do
@@
11:        Normalize batch loss: $\mathcal{L}_{\text{batch}}=\mathcal{L}_{\text{batch}}/|B_{k}|$.
@@
C.2.2 SFT of the Policy
Supervised fine-tuning (SFT) of the policy is performed on the dataset $D_{\text{SFT}}$ of prompts and target completions of the next reasoning step. The policy $\pi_{\theta}$ is instantiated as a general pretrained LLM. Specifically, we perform ’completion-only’ SFT such that for every (prompt, target completion) pair, the base model is trained to minimize the cross-entropy loss between its predicted token probabilities and the ground-truth target completion.
@@ C.2.3 Q-Value Model Training
The q-value model $Q_{\varphi}$ is trained on $D_{\text{QVM-train}}$ to assign appropriate scalar scores to the candidate reasoning trajectories. It is instantiated as a pre-trained LLM with additional linear layer and to which a shifted and rescaled classification head is added; we denote all of its trainable weights as $\varphi$. Depending on the reward design, the q-value model can be trained via scalar (least squares) regression if continuous rewards are chosen, or with a classification objective such as the Binary Cross-Entropy (BCE) loss, if trajectories are labelled with binary rewards or as chosen-rejected preference pairs.
By the end of training, $Q_{\varphi}$ should output accurate q-value scores, which will later guide policy refinement in Phase II and will improve the search accuracy when used in the MCTS.
Algorithm 3 Fine-Tuning the Q-Value Model $Q_{\varphi}$
Input: Q-value model $Q_{\varphi}$ (QVM), dataset $D_{\text{QVM-train}}=\{(x^{(i)},y^{(i)})\}$, training hyperparameters (optimizer, learning rate $\eta$, batch size $B$, and maximum epochs $E$).
@@
3:     Divide $D_{\text{QVM-train}}$ into batches $\{B_{k}\}$ of size $B$.
4:     for each batch $B_{k}$ do
5:        for each sample $(x^{(i)},y^{(i)})\in B_{k}$ do
@@
C.3 Training Phase 2: RL Tuning of Policy with MCTS
Phase 2 involves generating reasoning reasoning sequences from the policy with MCTS and the q-value model, and fine-tuning the policy with an RL-based alignment algorithm to generate better completions. The q-value model must also be continually updated in this training loop to keep in-distribution with the policy’s outputs. Sufficient Phase 1 pre-training of the policy and q-value model is crucial to ensure stable training of the models in Phase 2. The MCTS structure which provides a balanced exploration-exploitation search combined with repeated sampling of the policy ensures sufficient exploration during this online-RL phase. This final training phase returns the finetuned policy and q-value model.
C.3.1 Phase 2 Algorithm
Phase 2 uses a set $D_{p}=\{p^{(i)}\}$ of prompt questions - these questions may be isolated from the phase 1 dataset $D_{\text{SFT}}$. The training process (Algorithm 4) involves a repetition of a MCTS rollout phase followed by a training (reinforcement) phase.
Data generation: MCTS rollout. To obtain data for the training, a MCTS tree $\mathcal{T}^{(i)}$ is build w.r.t. each question $p^{(i)}$ using the algorithm in Algorithm  1 in training mode. The set of hyperparameters for MCTS $\Xi_{MCTS}$, denotes the number of MCTS iterations $N$ (per question), the number of children expanded in every selection phase $M$, exploration constants $c_{1},c_{2}$, and backpropagation weight $\alpha$. To enhance the quality of the data, we prune the generated MCTS tree $\tilde{\mathcal{T}}^{(i)}=(\tilde{N}^{(i)},\tilde{E}^{(i)})$ to only include paths that reached a terminal state since only these paths received the reward. Then, we extract all nodes and a set of node characteristics from the pruned tree. The dataset comprises of state, action and q-value triplets of the pruned tree: $\{(s_{j}^{(i)},z_{j}^{(i)},q(s_{j}^{(i)})\}_{s_{j}\in\tilde{N}^{(i)}}$ The data is stored in a replay buffer $\mathcal{R}$.
The training: RL phase. The reinforcement phase samples a batch of reasoning sequences from the replay buffer. From each trajectory, constituent states, actions and value estimatesand uses the corresponding values attributed during MCTS to perform RL training (for example with PPO or Reinforce). Alternative schemes may involve selecting preference pairs among trajectories and then aligning the policy using DPO, or simply selecting the most desirable trajectory per question and performing further SFT training.
During this reinforcement phase, the value model is updated to mimic the (backpropagated) values from the MCTS process (Algorithm 7).

### v4

@@ B.4.1 Evaluation Types
Absolute evaluations assign scalar values to each step, assessing aspects such as coherence, correctness, or helpfulness, using regression-based models. Moreover, evaluation dimensions can also be modeled as binary with classification models. While regression models provide more information, classification models capture correctness more naturally since a statement is usually correct or incorrect. On the other hand, the former ones are more suitable for measuring quality, such as the degree of coherence. Depending on the specific quality being evaluated, the choice between regression and classification models should align with the evaluation’s goals. Additionally, absolute scores can be transformed into rankings if needed, providing flexibility across various applications.
@@ B.4.2 Evaluation of Reasoning Steps
External model approaches include value models that predict scalar reward signals (reward models) [37, 155, 91], reinforcement learning values like state-values (v-value models) [138], state-action values (q-value models), or pairwise models like the Bradley-Terry and PairRM [75] frameworks. A more detailed comparison of reward models, v-value, and q-value models can be found in Appendix B.3.2.
There exist a large range of internal model approaches as substitutes for value models. They typically rely on methods like prompting the policy model to output scores. Examples include MCT Self-Refine (MCTSr) [181], querying for a binary feedback (e.g., “Is the answer correct? answer yes or no”) [184] and evaluating the probability of the output, leveraging uncertainty metrics such as token entropy or aggregated probabilities [187], and others [183].
@@ C.1.1 Setup and Notation
We will now present the details of the x1 training pipeline.
The MDP Design of x1 follows the definition presented in Appendix A.1 with the $\gamma$ values between $[0.95,1]$ to avoid over-penalizing long reasoning sequences. In the RLM setup, the state and action spaces of the underlying MDP constitute a tree in which every state $s$ other than the starting state $s_{0}$ has exactly one action $a_{s}$ leading to it. This allows us to simplify the notation by omitting actions wherever it’s clear from the context that we are referring only to an action leading to a given state. For every action $a$ leading from the state $s$ to the state $s^{\prime}$ we will write:
@@
$V_{\pi}(s_{t})$ $=\mathbb{E}_{\pi}\left[\gamma^{T-t}r(s_{T+1})\right]\quad\in[-1,1]$
@@
We use a pre-trained LM with parameters $\theta$ as a policy model, which we denote as $\pi_{\theta}$. The model autoregressively generates a sequence of tokens. We use a special token ’End of Intermediate Step’ (eois) to indicate the end of the reasoning step. We use a standard end-of-sequence (eos) token to indicate the end of the final reasoning step concluding the reasoning trajectory.
A parametric value model is used to evaluate the quality of states. While MCTS traditionally approximates these values through extensive simulations, such an approach is computationally expensive and impractical in the RLM context. Inspired by AlphaZero [138], which replaces simulations with a parameterized value model, we estimate the state-action values, i.e. the q-values, for reasoning sequences using a value model — effectively employing a process-based q-value model $Q_{\varphi}$ (see Appendix B.3). The value model is instantiated as a pre-trained transformer-based LM, modified by adding three linear layers and a shifted, rescaled sigmoid activation to align the output domain to the state action function domain $[-1,1]$ (see Eq. 2). This setup proved more stable than alternatives, such as a tanh activation or a cropped linear layer. We will show in the following how such a model can be trained and provide a description for the data generation process in Appendix D. During training, we assume access to a final answer verifier, which evaluates the correctness of the model’s final answer and provides the true reward.
@@ C.1.2 MCTS Algorithm
We now present the algorithmic steps of a Monte Carlo Tree Search variant similar to AlphaZero as implemented in the x1 reasoning framework, which we detail in Algorithm 1. The MCTS search operates in two distinct modes: training and inference. The core difference is that, during training, a final answer verifier evaluates and scores the final reasoning steps, providing a true reward signal that is backpropagated through the MCTS tree. This reward serves as a reliable learning signal for the value model $Q_{\varphi}$. During inference, however, the verifier is unavailable, and decisions rely solely on the value model.
Notation We chose to store all values in nodes instead of edges, which defines the following set of statistics saved for each node $s$:
  • $N(s)$ the visit count of node $s$,
@@
  • $\beta(s)$ the binary terminality function, returns $1$ if the node $s$ is terminal and 00 otherwise.
The Selection phase iteratively identifies the most promising child node with a selection policy, which in x1 is a node-based variant of the PUCT algorithm in AlphaZero [139] (which is defined on edge-based values) without a prior for selecting a child of $s$:
$\operatorname*{arg\,max}_{s_{c}\in\mathcal{C}(s)}\quad q(s_{c})+\frac{\sqrt{N(s)-1}}{1+N(s_{c})}\cdot\left(c_{1}+\log{\frac{N(s)+c_{2}}{c_{2}}}\right),$
@@
Expansion We append $M$ nodes to the selected leaf, $M$ being a hyperparameter. One of the major challenges in applying RLMs is maintaining the diversity of reasoning paths. By adding $M$ nodes, we increase the exploration of alternative reasoning paths.
The Backpropagation step serves to propagate information from the terminal nodes back to their ancestors. In our implementation, we update the running estimates of the q-values using the following formula:
@@
True Reward Propagation We improve the quality of the q-values by propagating the real final rewards back through the tree when a terminal state $s_{T+1}$ is reached. During training, terminal nodes can be evaluated against a reference golden answer $g^{*}$ using an external verifier. For actions leading to terminal states, the associated reward is equal to the q-value (see Eq. 2). Therefore, instead of using the prediction of the q-value model, we initialize $q(s_{T+1})$ with the true reward $r(s_{T+1})$ based on the evaluation of the external verifier. The reward is then backpropagated via the q-values through the tree with our backpropagation operator. This adjustment anchors the q-value model predictions with real reward signals and prevents the q-value model predictions to diverge.
Best Path Selection After $N$ iterations, MCTS will have formed a tree in which every path corresponds to one of the explored reasoning trajectories. The final reasoning step in a path with the highest terminal value estimate is returned as the final solution.
Algorithm 1 MCTS for Reasoning (Training mode in red)
@@
C.2 Training Phase 1
To adequately employ the MCTS-based reasoning scheme introduced in Appendix C.1, the policy model must be fine-tuned to generate responses in the format of semantically-relevant reasoning steps. The value model – a q-value model in our case – must be trained to accurately estimate the values of the sequences of reasoning steps.
The first phase focuses on preparing the policy and value models to generate and evaluate reasoning trajectories effectively. This is achieved by supervised fine-tuning (SFT) training on a dataset of example sequences of reasoning steps (where intermediate reasoning steps are terminated by an ”End of Intermediate Step” eois token). The objective is twofold: (1) to fine-tune the policy model $\pi_{\theta}$ to produce semantically coherent reasoning steps, and (2) to train the q-value model $Q_{\varphi}$ to accurately assign scalar scores to reasoning trajectories, distinguishing between high-quality and suboptimal reasoning paths.
This supervised fine-tuning phase ensures that the policy model can generate reasoning steps consistent with the structured format required for downstream MCTS-based exploration, while the q-value model provides reliable evaluations of intermediate and terminal states. Together, these components form the foundation for the subsequent online reinforcement learning in the second phase, where the policy and q-value models are further refined through interaction with the reasoning framework.
C.2.1 Datasets Generation and Preparation
Performing SFT of the policy model requires a dataset of high-quality reasoning sequences denoted as $D_{\text{SFT}}=\{\left(x_{\text{SFT}}^{(i)},\,y_{\text{SFT}}^{(i)}\right)\}$. Each pair in the dataset consists of a prompt $x_{\text{SFT}}^{(i)}$ composed of a sequence of reasoning steps (for example $x_{\text{SFT}}^{(i)}=(z_{0}^{(i)},...,z_{j}^{(i)})$), and a target completion $y_{\text{SFT}}^{(i)}=z_{j+1}^{(i)}$ which is the subsequent reasoning step or final answer. Appendix D contains a detailed account of the dataset creation and processing. It covers how the special eois token is appended to reasoning steps to mark the end of a step during inference.
Similarly to the policy model, training the q-value model requires a supervised dataset of reasoning sequences and corresponding scores. We denote this dataset $D_{\text{QVM-train}}=\{(x_{\text{QVM-train}}^{(i)},y_{\text{QVM-train}}^{(i)})\}$, with reasoning sequences $x_{\text{QVM-train}}^{(i)}=(z_{0}^{(i)},...,z_{t}^{(i)})$ and target q-value $y_{\text{QVM-train}}^{(i)}$. Appendix D explains how this dataset can be generated using an initial list of questions, a base LLM for querying, and a verifier program to label reasoning sequences as conducive to a correct final answer or not.
@@ C.2.1 Datasets Generation and Preparation
Input: Policy Model $\pi_{\theta}$, tokenized dataset $D_{\text{SFT}}=\{(x^{(i)},y^{(i)})\}$, training hyperparameters (optimizer, learning rate $\eta$, batch size $B$, and maximum number of epochs $E$).
@@
3:     Divide $D_{\text{SFT}}$ into batches $\{\mathcal{B}_{k}\}$ of size $B$.
4:     for each batch $\mathcal{B}_{k}$ do
@@
11:        Normalize batch loss: $\mathcal{L}_{\text{batch}}=\mathcal{L}_{\text{batch}}/|\mathcal{B}_{k}|$.
@@
C.2.2 SFT of the Policy Model
The supervised fine-tuning of the policy model, which we illustrate in Algorithm 2, is performed on the dataset $D_{\text{SFT}}$ of prompts and target completions of the next reasoning step. The policy $\pi_{\theta}$ is instantiated as a general pre-trained LLM. Specifically, we perform ’completion-only’ SFT such that for every (prompt, target completion) pair, the base model is trained to minimize the cross-entropy loss between its predicted token probabilities and the ground truth target completion.
@@ C.2.3 Q-Value Model Training
The q-value model $Q_{\varphi}$ is trained on $D_{\text{QVM-train}}$ to assign appropriate scalar scores to the candidate reasoning trajectories with the details being shown in Algorithm 3. It is instantiated as a pre-trained LLM with additional linear layers and to which a shifted and rescaled classification head is added; we denote all of its trainable weights as $\varphi$. Depending on the reward design, the q-value model can be trained via scalar (least squares) regression if continuous rewards are chosen, or with a classification objective such as the Binary Cross-Entropy (BCE) loss, if trajectories are labeled with binary rewards or as chosen-rejected preference pairs.
By the end of training, $Q_{\varphi}$ should output accurate q-value scores, which will later guide policy refinement in the second phase and improve the search accuracy when used in the MCTS.
Algorithm 3 Fine-tuning of the Q-Value Model $Q_{\varphi}$
Input: Q-value model $Q_{\varphi}$ (Q-VM), dataset $D_{\text{QVM-train}}=\{(x^{(i)},y^{(i)})\}$, training hyperparameters (optimizer, learning rate $\eta$, batch size $B$, and maximum number of epochs $E$).
@@
3:     Divide $D_{\text{QVM-train}}$ into batches $\{\mathcal{B}_{k}\}$ of size $B$.
4:     for each batch $\mathcal{B}_{k}$ do
5:        for each sample $(x^{(i)},y^{(i)})\in\mathcal{B}_{k}$ do
@@
C.3 Training Phase 2
Phase 2 involves generating reasoning sequences from the policy model with MCTS and the q-value model, and fine-tuning the policy model with an RL-based alignment algorithm to generate better completions. The q-value model must also be continually updated in this training loop to keep in-distribution with the policy model’s outputs. Sufficient pre-training of the policy and q-value models in the first phase is crucial to ensure stable training of these models in the second phase. The MCTS structure which provides a balanced exploration-exploitation search combined with repeated sampling of the policy model ensures sufficient exploration during this online-RL phase. This final training phase returns the fine-tuned policy and q-value models.
C.3.1 Training Algorithm
Phase 2 uses a set $D_{p}=\{p^{(i)}\}$ of prompt questions, which may be isolated from the phase 1 dataset $D_{\text{SFT}}$. The training process, displayed in Algorithm 4, involves a repetition of a MCTS rollout phase followed by a training (reinforcement) phase.
To obtain data for the training, a MCTS tree $\mathcal{T}^{(i)}$ is build w.r.t. each question $p^{(i)}$ using Algorithm 1 in training mode. The set of hyperparameters for MCTS $\Xi_{MCTS}$, denotes the number of MCTS iterations $N$ (per question), the number of children expanded in every selection phase $M$, exploration constants $c_{1},c_{2}$, and backpropagation weight $\alpha$. To enhance the quality of the data, we prune the generated MCTS tree $\tilde{\mathcal{T}}^{(i)}=(\tilde{N}^{(i)},\tilde{E}^{(i)})$ to only include paths that reached a terminal state since only these paths received the reward. Then, we extract all nodes and a set of node characteristics from the pruned tree. The dataset comprises of state, action and q-value triplets of the pruned tree: $\{(s_{j}^{(i)},z_{j}^{(i)},q(s_{j}^{(i)})\}_{s_{j}\in\tilde{N}^{(i)}}$. The data is stored in a replay buffer $\mathcal{R}$.
The reinforcement phase samples a batch of reasoning sequences from the replay buffer. From each trajectory, constituent states, actions, value estimates and the corresponding values attributed during MCTS are used to perform RL training (for example with PPO or reinforcement). Alternative schemes may involve selecting preference pairs among trajectories and then aligning the policy using DPO, or simply selecting the most desirable trajectory per question and performing further SFT training.
During this reinforcement phase, the value model is updated to mimic the (backpropagated) values from the MCTS process, which we illustrate in Algorithm 7.